An economist in Kenya argued that 3.5% of all SMEs would file for bank loans next
year. For a random sample of 100 SMEs, estimate the probability that at least three will
file bank loans next year. Assuming that the economist’s prediction is correct.
1
Expert's answer
2020-12-24T12:45:21-0500
"p_{100}^3=C_{100}^{3}\times 0.035^3\times (1-0.035)^{97}=0.22"